# Learning Hierarchical Sparse Transform Coding of 3DGS

### v2

@@ 5.2.2 Evaluating Computational Cost
In terms of parameter overhead, the transforms in SHTC introduce 5,093 additional parameters over the HAC baseline. However, when SHTC is integrated into HAC, the entropy model queries the hash grids and feeds an MLP that predicts probability parameters for low-dimensional transform coefficients instead of high-dimensional anchor features. This change reduces the size of this MLP. As a result, the integration of SHTC into HAC incurs a net overhead of only 1,154 parameters. In contrast, HAC++ increases the number of parameters by 45,400 relative to HAC. These results highlight the parameter efficiency of SHTC, which stems from a sparsity-aware transform design rather than simply adopting a larger black-box MLP.
@@ 5.2.3 Visual Comparison
Due to space limitations, visual comparisons are provided in the supplementary material.
(a) Mip-NeRF360
(b) Tank&Temples
(c) DeepBlending
@@ 5.3 Comparison with Other Methods
In addition to the comparisons with anchor-based compression methods using R–D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines [kerbl20233d, lu2024scaffold, lee2024compact, niedermayr2024compressed, girish2024eagles, fan2024lightgaussian, morgenstern2024compact, navaneet2024compgs, papantonakis2024reducing, wang2024end, liu2024compgs, chen2025fast, shin2025localityaware, tang2025neuralgs, hanson2025pup, xie2025sizegs, tian2025flexgaussian, lee2025optimized, tang2025feature, ma2025enhancing, morgenstern2024compact, lee2025compression, xie2024mesongs, yang2025hybridgs]. Because many of these methods only report one or two operating points on the R-D curve, they do not support a meaningful BD-rate computation; for these methods, we therefore provide a numerical comparison in tabular form in the supplementary material.
[…]
@@ 5.4 Ablation Study
As shown in Fig. 3, using only the base layer of SHTC, denoted as ‘Ours SHTC (BL)’, achieves slightly better or comparable R–D performance compared to HAC++ at low and medium bitrates. However, the improvements are modest overall, and SHTC (BL) fails to consistently surpass HAC++ in the high-rate region. This limited gain is mainly due to truncation-induced information loss in the base layer, which inevitably discards some useful information and constrains the achievable performance. In contrast, when we use the full SHTC to compress anchor attributes, which includes an additional neural refinement layer to compensate for truncation-induced information loss, the proposed method yields consistent gains over HAC++ across both low- and high-rate regions. At high bitrates, it also remains clearly superior in terms of R-D performance and does not suffer from the high-rate quality bottleneck observed with the base-layer-only variant.
The supplementary material provides a more detailed quantitative ablation study and a further analysis of the choice of transforms, and also discusses the limitations of our approach and directions for future work.
@@ 6 Conclusion
In this paper, we introduced SHTC, to our knowledge, the first end-to-end optimized transform coding framework for 3DGS compression, which achieves substantial R–D gains and improved inference efficiency with only minimal parameter overhead. By leveraging learned analysis–synthesis transforms for decorrelation instead of relying solely on highly complex entropy models, SHTC improves compression efficiency and reduces decoding latency, making it well suited to consumer-facing scenarios where fast transmission and decoding are critical. SHTC has the potential to inspire future work on 3DGS compression, for example by motivating more expressive transform architectures and spatially adaptive variants. More broadly, the idea of incorporating signal priors to improve the parameter efficiency of neural codecs may benefit the design of low-complexity neural image and video codecs.
[…]
Supplementary Material
(a) Synthetic-NeRF
(b) BungeeNeRF
Figure 5: Comparison of our method with anchor-based methods on two additional datasets, Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf].
Figure 6: Visual comparison between HAC++ and our SHTC method on four scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), ‘garden’ and ‘stump’ (Mip-NeRF360).
7 Overview of the Supplementary Material
This supplementary document provides additional technical details, extended experiments, and ablation studies that complement the main paper. Its content is organized as follows:
  • Sec. 8 compares our method with representative anchor-based methods on two additional datasets, Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf].
  • Sec. 9 presents additional comparisons of computational cost between our approach and CAT-3DGS.
  • Sec. 10 shows visual comparisons.
  • Sec. 11 presents more detailed ablation studies.
  • Sec. 12 provides a more detailed discussion of our transform choices.
  • Sec. 13 presents additional comparisons in the training time.
  • Sec. 14 provides further comparisons with a set of widely used baselines and recent methods.
  • Sec. 15 summarizes implementation details.
    • Sec. 15.1 explains how we compute pixel-wise distortion in the YCbCr color space.
    • Sec. 15.2 details the selection of the $\lambda$ values used to generate the R-D curves in the main paper.
  • Sec. 16 discusses the limitations of our approach and outlines directions for future work.
  • Sec. 17 reviews additional related work to further contextualize our method.
8 Additional Results on the Synthetic-NeRF and BungeeNeRF Datasets
Fig. 5 compares the R-D curves of our method with several anchor-based 3DGS compression methods on the Synthetic-NeRF [mildenhall2021nerf] and BungeeNeRF [xiangli2022bungeenerf] datasets. As shown in the figure, our method consistently achieves better R-D performance, attaining a smaller memory footprint at comparable PSNR levels. Tab. 3 further reports the BD-rate gains of our method over three anchor-based baselines. Taken together, these additional experimental results further demonstrate the effectiveness of our approach and serve as a supplementary comparison to the experiments reported in the main text.
Dataset HAC++ CAT-3DGS (40k) HAC
Synthetic-NeRF -13.45% - -26.33%
BungeeNeRF -10.04% -39.90% -52.05%
Table 3: The BD-rate of our method relative to three baselines: HAC++ [chen2025hac++], CAT-3DGS [zhan2025catdgs] and HAC [chen2024hac]. Negative values indicate bitrate savings (better) compared with each baseline.
9 Additional Comparisons with CAT-3DGS
[…]
Table 4: Evaluating the effectiveness: CAT-3DGS vs. SHTC + simple entropy model. From left to right, we report four metrics: encoding time, decoding time, and BD-rate relative to the HAC baseline. The best result for each metric is highlighted in bold.
As a supplement to Table 2 in the main paper, Tab. 4 compares our method with CAT-3DGS. The evaluation metrics include encoding time, decoding time, and BD-rate relative to the HAC baseline. With HAC as the common baseline, our method achieves a larger reduction in memory footprint while requiring substantially less encoding and decoding time at inference. For CAT-3DGS [zhan2025catdgs], we encountered out-of-memory (OOM) errors during rendering when running the official implementation on our hardware, so the rendering FPS could not be measured. However, the training time, encoding time, and decoding time can still be measured before this OOM issue occurs. Consequently, we omit the rendering-FPS comparison for CAT-3DGS in the table.
10 Visual Comparison
@@
We provide visual comparisons between our SHTC approach and HAC++ [chen2025hac++] in Fig. 6. For each example, we report the bitstream size (in MB) and the PSNR (in dB) of the shown crop. To illustrate how SHTC reduces rendering distortion under a smaller or comparable memory footprint, we evaluate four representative scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), and ‘garden’ and ‘stump’ (Mip-NeRF360), which are shown from the first to the fourth row in Fig. 6. In the ‘playroom’ scene, our method better preserves the structure of the wall-mounted switch, which appears as an unrecognizable blur in the HAC++ reconstruction. In the ‘train’ scene, our method suppresses shading and dark banding artifacts along the painted stripes. In the ‘garden’ scene, our method removes the black floater near the table edge. In the ‘stump’ scene, our method effectively reduces floater artifacts around the tree stump.
@@
11 Additional Ablation Study
11.1 Additional Quantitative Ablation Study
@@
We plot the R–D curves of these variants on the Mip-NeRF360 dataset in Fig. 7(a). Mip-NeRF360 contains diverse and challenging indoor and outdoor scenes, making it a suitable benchmark for ablation with reduced bias. From Fig. 7(a), we observe that introducing a KLT-based base layer for anchor feature compression (HAC-E + BL(F15)) already yields a clear gain over the HAC-E baseline and brings the performance close to HAC++. Extending the base layer to also compress anchor scalings (HAC-E + BL(F15+S)) and further switching to the YCbCr-based distortion metric (HAC-E + BL(F15+S) + YCbCr) provide additional, consistent improvements. The complete SHTC configuration trained with the YCbCr-based distortion (HAC-E + Full SHTC + YCbCr) achieves the best overall R-D trade-off among all variants.
Tab. 5 quantifies these effects by reporting BD-rate values over HAC-E on Mip-NeRF360. HAC++ is also included to facilitate a direct comparison between the two paradigms: using learned transforms with a simple entropy model versus relying solely on a highly complex entropy model. The results show that even the simplest transform-based variant (HAC-E + BL(F15)) already surpasses HAC++, and each additional component yields further BD-rate reductions, leading to a substantial $29.33\%$ reduction for HAC-E + Full SHTC + YCbCr.
@@
11.2 Comparison with No-Truncation Variants
Finally, we investigate whether directly transmitting all KLT coefficients can serve as a competitive alternative to our truncated-plus-refinement design. In Fig. 7(b), we first consider a variant that does not truncate the KLT coefficients of anchor features, denoted by HAC-E + BL(F50+S) + YCbCr. This configuration directly quantizes and entropy-codes all 50 KLT coefficients. Under the same rate-control parameter $\lambda$, it attains higher PSNR than the truncated-only baseline, but at a significantly higher bitrate, so it offers no improvement in overall R-D performance compared with simply truncating the coefficients.
@@ 11.2 Comparison with No-Truncation Variants
$q_{s}^{(i)}=q_{s}\exp(\alpha i),\qquad i\in\{0,\ldots,49\},$ (10)
@@
12 Discussion of Transform Choices
12.1 Transform Choices for the Base Layer
@@
We considered several candidates for the base-layer transform, including the classical DCT and the Haar wavelet. To guide this choice, we compared their ability to decorrelate anchor channels and compact energy. Fig. 8 shows the absolute inter-channel correlation matrices and the normalized energy distributions of the original HAC features, the DCT coefficients, the Haar coefficients, and the KLT coefficients. DCT and Haar can reduce inter-channel correlation and slightly improve energy compactness, but their gains are limited: the correlation matrices still exhibit noticeable off-diagonal structure and the energy remains relatively spread out across channels. This behavior is expected because their bases are fixed and data-agnostic. In contrast, KLT yields an almost perfectly diagonal correlation matrix and concentrates most of the energy in the first few coefficients. This data-dependent transform is obtained as the eigenvectors of the sample covariance and is theoretically optimal for decorrelation and energy compaction among linear transforms. The main additional cost, relative to DCT and Haar, is storing the KLT basis. Even if the full basis must be transmitted, a 50-dimensional feature vector only requires a $50\times 50$ matrix (2,500 coefficients), which is negligible compared with the overall bitstream size.
@@
Beyond classical transforms, we also compare our ‘HAC-E + KLT (F15)’ variant (the simplest version of SHTC) with a variant equipped with an MLP-based analysis and synthesis transform pair. Both networks are 5-layer MLPs, with a total of 21,930 parameters that must be transmitted as part of the bitstream. This comparison, evaluated on the Mip-NeRF360 dataset and shown in Fig. 9, reveals that the black-box MLP transform performs significantly worse in terms of the R-D trade-off than the simple KLT-based design. A likely reason is that, under a strict parameter budget, the MLP-based transform lacks sufficient expressive power to accurately reconstruct the input, leading to larger reconstruction errors and thus worse R-D performance.
@@
Figure 9: Comparison between ‘HAC-E + KLT (F15)’ and ‘HAC-E + MLP’ on the Mip-NeRF360 dataset.
12.2 Transform Choices for the Refinement Layer
@@
To verify this intuition, we conduct several offline experiments; here we present results on the bicycle scene of the Mip-NeRF360 dataset, where we compress the KLT residual in two different ways. The first variant uses a pure MLP-based transform with an MLP analysis and synthesis transform pair; both networks are 5-layer MLPs, resulting in a total of 21,930 parameters that must be transmitted in the bitstream. The second variant uses our sparsity-aware refinement layer, which contains only 5,093 parameters. We train both models for 3,000 iterations and measure the $\ell_{1}$ reconstruction loss of the KLT residual. As shown in Fig. 10, the sparsity-aware transform consistently drives the loss down to a much lower level than the MLP-based transform, despite using roughly four times fewer parameters. This early-stage verification supports our motivation: leveraging the sparsity prior and using a compressed-sensing-inspired architecture for residual coding yields a highly parameter-efficient refinement module. For this reason, we adopt the sparsity-aware design as our default choice for the neural refinement layer in SHTC.
Figure 10: Offline comparison of our sparsity-aware refinement layer and a black-box MLP-based transform on the bicycle scene of the Mip-NeRF360 dataset. The plot shows the $\ell_{1}$ reconstruction loss of the KLT residual versus training iteration.
13 Training time
Tab. 6 compares the training time of our method with three advanced anchor-based 3DGS compression methods that rely on complex entropy models. Although our method is not the fastest to train, its training time remains moderate: it requires more training time than HAC++, but consistently less than CAT-3DGS and ContextGS across all three benchmarks. Designing a compression system that simultaneously attains the best possible training efficiency, R-D performance, and inference speed is notoriously difficult; in practice, there is no free lunch. We argue that the resulting overhead in training time is acceptable because 3DGS compression follows a “compress once, decompress many times” paradigm. Training is a one-time offline cost borne by 3DGS asset producers, whereas decoding is performed repeatedly on the user side. End users are therefore insensitive to training time but directly affected by download and decoding latency. In this context, trading a modest increase in training time for improved R-D performance and faster decoding is both realistic and, in practice, a favorable design choice. Our method realizes this trade-off in practice, achieving improved R-D performance and lower decoding latency while keeping the training time within a reasonable range.
CAT-3DGS ContextGS HAC++ Ours
Mip-NeRF360 5725 3927 2735 3218
Tank&Temples 4034 2505 1880 2315
DeepBlending 3353 2320 1857 2270
Table 6: Training time (in seconds) of CAT-3DGS, ContextGS, HAC++, and our method on three datasets. The best and second-best results are highlighted in red and yellow cells, respectively.
Datasets Mip-NeRF360 Tank&Temples DeepBlending
@@ 13 Training time
3DGS [kerbl20233d] 27.46 0.812 0.222 750.9 23.69 0.844 0.178 431.0 29.42 0.899 0.247 663.9
Scaffold-GS [lu2024scaffold] 27.50 0.806 0.252 253.9 23.96 0.853 0.177 86.50 30.21 0.906 0.254 66.00
Compact3DGS [lee2024compact] 27.08 0.798 0.247 48.80 23.32 0.831 0.201 39.43 29.79 0.901 0.258 43.21
Compressed3D [niedermayr2024compressed] 26.98 0.801 0.238 28.80 23.32 0.832 0.194 17.28 29.38 0.898 0.253 25.30
EAGLES [girish2024eagles] 27.14 0.809 0.231 58.91 23.28 0.835 0.203 28.99 29.72 0.906 0.249 52.34
LightGaussian [fan2024lightgaussian] 27.00 0.799 0.249 44.54 22.83 0.822 0.242 22.43 27.01 0.872 0.308 33.94
Navaneet et al. [navaneet2024compgs] 27.12 0.806 0.240 19.33 23.44 0.838 0.198 12.50 29.90 0.907 0.251 13.50
Reduced3DGS [papantonakis2024reducing] 27.19 0.807 0.230 29.54 23.57 0.840 0.188 14.00 29.63 0.902 0.249 18.00
RDOGaussian [wang2024end] 27.05 0.802 0.239 23.46 23.34 0.835 0.195 12.03 29.63 0.902 0.252 18.00
PUP 3D-GS [hanson2025pup] 26.67 0.786 0.272 74.65 22.72 0.801 0.244 43.33 28.85 0.881 0.302 69.92
SizeGS [xie2025sizegs] 27.48 0.806 0.240 18.17 24.04 0.840 0.200 10.93 30.24 0.903 0.271 7.92
FlexGaussian [tian2025flexgaussian] 26.38 0.780 0.251 40.80 22.44 0.804 0.219 16.30 28.61 0.884 0.269 25.48
NeuralGS [tang2025neuralgs] 27.35 0.806 0.240 16.90 23.63 0.841 0.192 12.06 29.91 0.906 0.254 12.98
LocoGS [shin2025localityaware] 27.40 0.815 0.219 13.89 23.89 0.854 0.160 12.34 30.17 0.906 0.244 13.39
FCGS (low rate) [chen2025fast] 27.05 0.798 0.237 36.30 23.48 0.833 0.193 18.80 29.27 0.893 0.257 30.10
FCGS (high rate) [chen2025fast] 27.39 0.806 0.226 67.20 23.62 0.839 0.184 33.60 29.58 0.899 0.248 54.50
SOG [morgenstern2024compact] 26.56 0.791 0.241 16.70 23.15 0.828 0.198 9.30 29.12 0.892 0.270 5.70
CodecGS [lee2025compression] 27.30 0.810 0.236 9.78 23.63 0.842 0.192 7.46 29.82 0.907 0.251 8.62
MesonGS [xie2024mesongs] 26.99 0.796 0.247 27.16 23.32 0.837 0.193 16.99 29.51 0.901 0.251 24.76
CompGS [liu2024compgs] 27.26 0.803 0.239 16.50 23.70 0.837 0.208 9.60 29.69 0.901 0.279 8.77
FP-Net [ma2025enhancing] 27.46 0.801 0.249 11.08 24.13 0.847 0.192 6.34 30.14 0.906 0.274 3.46
FP-GS [tang2025feature] 27.53 0.806 0.240 12.82 24.32 0.811 0.185 7.32 30.05 0.906 0.266 3.91
HAC [chen2024hac] 27.53 0.807 0.238 15.26 24.04 0.846 0.187 8.10 29.98 0.902 0.269 4.35
Context-GS [wang2024contextgs] 27.62 0.808 0.237 12.68 24.20 0.852 0.184 7.05 30.11 0.907 0.258 3.45
HAC++ [chen2025hac++] 27.60 0.803 0.253 8.34 24.22 0.849 0.190 5.18 30.16 0.907 0.266 2.91
Ours 27.69 0.804 0.248 8.10 24.34 0.845 0.191 5.19 30.19 0.903 0.272 2.49
14 Additional Comparisons with Other Methods

### v3

@@ 4.2.2 Evaluating Transform Overhead
The transforms in SHTC introduce 5,093 additional parameters over the HAC baseline. However, when SHTC is integrated into HAC, the entropy model queries the hash grids and feeds an MLP that predicts probability parameters for low-dimensional transform coefficients instead of high-dimensional anchor features. This change reduces the size of this MLP. As a result, the integration of SHTC into HAC incurs a net overhead of only 1,154 parameters. In contrast, HAC++ increases the number of parameters by 45,400 relative to HAC. These results highlight the parameter efficiency of SHTC, which stems from the sparsity-aware transform design rather than simply adopting a larger black-box MLP.
@@ 4.2.3 Visual Comparison
Due to space limitations, visual comparisons are provided in Appendix A.
@@ 4.3 Comparison with Other Methods
In addition to the comparisons with anchor-based compression methods using R-D curves and BD-rate, we also evaluate our method against a broad set of anchor-free and hybrid baselines. Because many of these methods only report one or two operating points on the R-D curve, they do not support a meaningful BD-rate computation; for these methods, we therefore provide a numerical comparison in tabular form in Appendix D and Table 4 in Appendix.
In addition, we provide a dedicated comparison with several conceptually representative baselines: the classic 3DGS (Kerbl et al., 2023) and Scaffold-GS (Lu et al., 2024), the current state-of-the-art (SOTA) anchor-free compression method OMG (NeurIPS 2025) (Lee et al., 2025a), the feed-forward compression method FCGS (Chen et al., 2025b), and four representative methods that apply transform coding in a post-training stage for 3DGS compression: SOG and MesonGS (ECCV 2024) (Morgenstern et al., 2024; Xie et al., 2024), CodecGS (ICCV 2025) (Lee et al., 2025b), HybridGS (ICML 2025) (Yang et al., 2025). Note that FCGS can also be seen as a method that applies transform coding in a post-training stage after 3DGS is fixed.
As shown in Fig. 13 in Appendix, our method consistently outperforms the above eight methods. These results clearly demonstrate that jointly optimizing the GS representation, entropy models, and the transforms in an end-to-end manner is substantially more effective than the ad hoc post-training transform paradigm where 3DGS and the transforms cannot be mutually adapted.
@@ 4.4 Ablation Study
As shown in Figure 6, using only the base layer of SHTC, denoted as ‘Ours SHTC (BL)’, achieves slightly better or comparable R–D performance compared to HAC++ at low and medium bitrates. However, the improvements are modest overall, and SHTC (BL) fails to consistently surpass HAC++ in the high-rate region. This limited gain is mainly due to truncation-induced information loss in the base layer, which inevitably discards some useful information and constrains the achievable performance. In contrast, when we use the full SHTC to compress anchor attributes, which includes an additional neural refinement layer to compensate for truncation-induced information loss, the proposed method yields consistent gains over HAC++ across both low- and high-rate regions. At high bitrates, it also remains clearly superior in terms of R-D performance and does not suffer from the high-rate quality bottleneck observed with the base-layer-only variant.
We conduct a thorough ablation study to isolate the contribution of each component; the main results are reported in Fig. 10(a) and Table 3 in Appendix, with additional details deferred to Appendix B.1. In addition to component-wise effects, a natural question is whether residual transform coding is actually necessary for handling truncation. To answer this, we compare (i) our residual transform coding scheme that compensates truncation error against (ii) a direct baseline that transmits all KLT coefficients without truncation. This comparison is provided in Appendix B.2. Finally, since SHTC introduces a new transform design, one may wonder why we adopt this particular design instead of more standard alternatives (e.g., DCT, DWT, or a generic MLP). We therefore include a dedicated analysis of these transform choices in Appendix C.
Limitations and future work are discussed in Appendix F.
@@ 5 Conclusion
In this paper, we introduce TTC, a new paradigm for 3DGS compression that jointly optimizes the 3DGS representation, entropy model, and analysis–synthesis transform under a unified R–D objective. Compared with anchor-based compressors that solely rely on increasingly complex entropy models and PTC methods that cannot co-adapt the transform with 3DGS, TTC provides a more effective route to redundancy removal. Within TTC, we propose SHTC, achieving substantial R–D gains and improved inference efficiency with minimal parameter overhead. More broadly, our parameter-efficient design may benefit low-complexity neural image and video codec design.
Impact Statement
Our method serves purely as a compression tool, which operates solely on existing data to reduce its storage or transmission cost. It neither generates new data nor edits the original content, thereby minimizing risks of misuse, such as data fabrication or manipulation. Given its functional scope, we do not anticipate any significant negative societal impacts associated with the deployment of our method.
On the positive side, by significantly reducing the file size of 3DGS representations and lowering decoding latency, our method facilitates faster downloading and decoding. This improves user experience in bandwidth-constrained or real-time applications such as virtual reality, immersive gaming, architectural visualization, and cultural heritage preservation.
Appendix A Visual Comparison
@@
We provide visual comparisons between our SHTC approach and HAC++ (Chen et al., 2025c) in Figure 9. For each example, we report the bitstream size (in MB) and the PSNR (in dB) of the shown crop. To illustrate how SHTC reduces rendering distortion under a smaller or comparable memory footprint, we evaluate four representative scenes: ‘playroom’ (DeepBlending), ‘train’ (Tanks & Temples), and ‘garden’ and ‘stump’ (Mip-NeRF360), which are shown from the first to the fourth row in Figure 9. In the ‘playroom’ scene, our method better preserves the structure of the wall-mounted switch, which appears as an unrecognizable blur in the HAC++ reconstruction. In the ‘train’ scene, our method suppresses shading and dark banding artifacts along the painted stripes. In the ‘garden’ scene, our method removes the black floater near the table edge. In the ‘stump’ scene, our method effectively reduces floater artifacts around the tree stump.
@@
Appendix B Additional Ablation Study
B.1 Additional Quantitative Ablation Study on Component Contributions
@@
We plot the R–D curves of these variants on the Mip-NeRF360 dataset in Figure 10(a). Mip-NeRF360 contains diverse and challenging indoor and outdoor scenes, making it a suitable benchmark for ablation with reduced bias. From Figure 10(a), we observe that introducing a KLT-based base layer for anchor feature compression (HAC-E + BL(F15)) already yields a clear gain over the HAC-E baseline and brings the performance close to HAC++. Extending the base layer to also compress anchor scalings (HAC-E + BL(F15+S)) and further switching to the YCbCr-based distortion metric (HAC-E + BL(F15+S) + YCbCr) provide additional, consistent improvements. The complete SHTC configuration trained with the YCbCr-based distortion (HAC-E + Full SHTC + YCbCr) achieves the best overall R-D trade-off among all variants.
Table 3 quantifies these effects by reporting BD-rate values over HAC-E on Mip-NeRF360. HAC++ is also included to facilitate a direct comparison between the two paradigms: using learned transforms with a simple entropy model versus relying solely on a highly complex entropy model. The results show that even the simplest transform-based variant (HAC-E + BL(F15)) already surpasses HAC++, and each additional component yields further BD-rate reductions, leading to a substantial $29.33\%$ reduction for HAC-E + Full SHTC + YCbCr.
@@
B.2 Comparison with No-Truncation Variants
Finally, we investigate whether directly transmitting all KLT coefficients can serve as a competitive alternative to our truncated-plus-refinement design. In Figure 10(b), we first consider a variant that does not truncate the KLT coefficients of anchor features, denoted by HAC-E + BL(F50+S) + YCbCr. This configuration directly quantizes and entropy-codes all 50 KLT coefficients. Under the same rate-control parameter $\lambda$, it attains higher PSNR than the truncated-only baseline, but at a significantly higher bitrate, so it offers no improvement in overall R-D performance compared with simply truncating the coefficients.
@@ B.2 Comparison with No-Truncation Variants
$q_{s}^{(i)}=q_{s}\exp(\alpha i),\quad i\in\{0,\ldots,49\},$ (8)
@@
Appendix C Discussion of Transform Choices
C.1 Transform Choices for the Base Layer
@@
We considered several candidates for the base-layer transform, including the classical DCT and the Haar wavelet. To guide this choice, we compared their ability to decorrelate anchor channels and compact energy. Figure 11 shows the absolute inter-channel correlation matrices and the normalized energy distributions of the original HAC features, the DCT coefficients, the Haar coefficients, and the KLT coefficients. DCT and Haar can reduce inter-channel correlation and slightly improve energy compactness, but their gains are limited: the correlation matrices still exhibit noticeable off-diagonal structure and the energy remains relatively spread out across channels. This behavior is expected because their bases are fixed and data-agnostic. In contrast, KLT yields an almost perfectly diagonal correlation matrix and concentrates most of the energy in the first few coefficients. This data-dependent transform is obtained as the eigenvectors of the sample covariance and is theoretically optimal for decorrelation and energy compaction among linear transforms. The main additional cost, relative to DCT and Haar, is storing the KLT basis. Even if the full basis must be transmitted, a 50-dimensional feature vector only requires a $50\times 50$ matrix (2,500 coefficients), which is negligible compared with the overall bitstream size.
@@
Beyond classical transforms, we also compare our ‘HAC-E + KLT (F15)’ variant (the simplest version of SHTC) with a variant equipped with an MLP-based analysis and synthesis transform pair. Both networks are 5-layer MLPs, with a total of 21,930 parameters that must be transmitted as part of the bitstream. This comparison, evaluated on the Mip-NeRF360 dataset and shown in Figure 12(a), reveals that the black-box MLP transform performs significantly worse in terms of the R-D trade-off than the simple KLT-based design. A likely reason is that, under a strict parameter budget, the MLP-based transform lacks sufficient expressive power to accurately reconstruct the input, leading to larger reconstruction errors and thus worse R-D performance.
@@
(a) Comparison between ‘HAC-E + KLT (F15)’ and ‘HAC-E + MLP’ on the Mip-NeRF360 dataset.
(b) Offline comparison of our sparsity-aware refinement layer and a black-box MLP-based transform on the bicycle scene of the Mip-NeRF360 dataset. The plot shows the $\ell_{1}$ reconstruction loss of the KLT residual versus training iteration.
Figure 12: Comparison of different design choices.
C.2 Transform Choices for the Refinement Layer
@@
To verify this intuition, we conduct several offline experiments; here we present results on the bicycle scene of the Mip-NeRF360 dataset, where we compress the KLT residual in two different ways. The first variant uses a pure MLP-based transform with an MLP analysis and synthesis transform pair; both networks are 5-layer MLPs, resulting in a total of 21,930 parameters that must be transmitted in the bitstream. The second variant uses our sparsity-aware refinement layer, which contains only 5,093 parameters. We train both models for 3,000 iterations and measure the $\ell_{1}$ reconstruction loss of the KLT residual. As shown in Figure 12(b), the sparsity-aware transform consistently drives the loss down to a much lower level than the MLP-based transform, despite using roughly four times fewer parameters. This early-stage verification supports our motivation: leveraging the sparsity prior and using a compressed-sensing-inspired architecture for residual coding yields a highly parameter-efficient refinement module. For this reason, we adopt the sparsity-aware design as our default choice for the neural refinement layer in SHTC.
@@
(a) Mip-NeRF360
(b) Tank&Temples
(c) DeepBlending
Appendix D Additional Comparisons with Other Methods